# Architect of the Bits World: Masked Autoregressive Modeling for Circuit Generation Guided by Truth Table

### v1

@@ Abstract
Logic synthesis, a critical stage in electronic design automation (EDA), optimizes gate-level circuits to minimize power consumption and area occupancy in integrated circuits (ICs). Traditional logic synthesis tools rely on human-designed heuristics, often yielding suboptimal results. Although differentiable architecture search (DAS) has shown promise in generating circuits from truth tables, it faces challenges such as high computational complexity, convergence to local optima, and extensive hyperparameter tuning. Consequently, we propose a novel approach integrating conditional generative models with DAS for circuit generation. Our approach first introduces CircuitVQ, a circuit tokenizer trained based on our Circuit AutoEncoder. We then develop CircuitAR, a masked autoregressive model leveraging CircuitVQ as the tokenizer. CircuitAR can generate preliminary circuit structures from truth tables, which guide DAS in producing functionally equivalent circuits. Notably, we observe the scalability and emergent capability in generating complex circuit structures of our CircuitAR models. Extensive experiments also show the superior performance of our method. This research bridges the gap between probabilistic generative models and precise circuit generation, offering a robust solution for logic synthesis.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Illustration of differentiable CircuitNN. CircuitNN is designed based on differentiable NAND gates. After DAS is guided by PI and PO pairs of the truth table, CircuitNN can get the precise circuit architecture logic equivalent to the truth table.
@@
Logic synthesis (Hachtel & Somenzi, 2005), a critical step in electronic design automation (EDA), transforms behavioral-level circuit designs into optimized gate-level circuits, ultimately yielding the final IC layout. The primary goal of logic synthesis is to identify the physical implementation with the fewest gates for a given circuit function. This task constitutes a challenging NP-hard combinatorial optimization problem. Current logic synthesis tools (Brayton & Mishchenko, 2010; Wolf et al., 2013) rely on human-designed heuristics, often leading to sub-optimal outcomes.
@@
Differentiable architecture search (DAS) techniques (Liu et al., 2018; Chu et al., 2020) offer novel perspectives on addressing challenges in this problem. Circuit functions can be represented through truth tables, which map binary inputs to their corresponding outputs. Truth tables provide a precise representation of input-output relationships, ensuring the design of functionally equivalent circuits. Inspired by this, researchers (Google Deepmind, 2023; Wang et al., 2024) have begun exploring the application of DAS to synthesize circuits directly from truth tables. Specifically, Google Deepmind (2023) proposed CircuitNN, a framework that learns differentiable connection structures with logic gates, enabling the automatic generation of logic circuits from truth tables. This approach significantly reduces the complexity of traditional circuit generation. Building on this, Wang et al. (2024) introduced T-Net, a triangle-shaped variant of CircuitNN, incorporating regularization techniques to enhance the efficiency of DAS.
@@
Despite these advancements, several challenges remain. The computational complexity of DAS grows quadratically with the number of gates, posing scalability issues. Although triangle-shaped architecture (Wang et al., 2024) partially mitigates this problem, redundancy persists. Additionally, DAS is susceptible to converging to local optima (Liu et al., 2018) and hyperparameters (network depth and layer width) require extensive searches. The challenges arise from the vast search space in DAS. Intuitively, limiting the search space through predefined parameters (network depth, gates per layer, and connection probabilities) can significantly reduce the complexity.
@@
Recent advances (OpenAI, 2023; Abramson et al., 2024; Esser et al., 2024; Li et al., 2024a) in conditional generative models have demonstrated remarkable performance across language, vision, and graph generation tasks. Motivated by these developments, we propose a novel approach to circuit generation that generates preliminary circuit structures to guide DAS in generating refined circuits matching specified truth tables. Firstly, we introduce CircuitVQ, a tokenizer with a discrete codebook for circuit tokenization. Built upon our Circuit AutoEncoder framework (Hou et al., 2022; Li et al., 2023a; Wu et al., 2025), CircuitVQ is trained through a circuit reconstruction task. Specifically, the CircuitVQ encoder encodes input circuits into discrete tokens using a learnable codebook, while the decoder reconstructs the circuit adjacency matrix based on these tokens. Subsequently, the CircuitVQ encoder serves as a circuit tokenizer for CircuitAR pretraining, which employs a masked autoregressive modeling paradigm (Chang et al., 2022; Li et al., 2023b). In this process, the discrete codes function as supervision signals. After training, CircuitAR can generate discrete tokens progressively, which can be decoded into initial circuit structures by the decoder of the CircuitVQ. These prior insights can guide DAS in producing refined circuits that match the target truth tables precisely.
Our key contributions can be summarized as follows:
@@ 2.1 Modeling Circuit as DAG
In this work, we model the circuit as a directed acyclic graph (DAG) (Brummayer & Biere, 2006), which facilitates graph autoregressive modeling. Specifically, each node in the DAG corresponds to a logic gate, such as AND, OR, NAND, or other gate types, while the directed edges represent the connections between these components.
@@ 2.2 Differentiable CircuitNN
As depicted in Figure 1, CircuitNN (Google Deepmind, 2023) replaces traditional neural network layers with logic gates (e.g., NAND) as basic computational units, learning to synthesize circuits by optimizing logic correctness based on truth tables. During training, input connections of each gate are determined through learnable probability distributions, enabling adaptive circuit architecture modification. To enable gradient-based learning, CircuitNN transforms discrete logic operations into continuous, differentiable functions using NAND gates for simplicity. The NAND gate is logically complete, allowing the construction of any complex logic circuit. Its continuous relaxation can be defined as:
@@
Additionally, CircuitNN employs Gumbel-Softmax (Jang et al., 2016) for stochastic sampling of gate inputs. Through stochastic relaxation, gate and network outputs are no longer binary but take continuous values ranging from 0 to 1 instead. This end-to-end differentiability allows the model to learn gate input distributions using gradient descent. After training, the continuous, probabilistic circuit is converted back into a discrete logic circuit by selecting the most probable connections based on the learned probability distributions as shown in Figure 1.
Figure 2: The training process of CircuitVQ, designed based on the circuit autoencoder and a vector quantizer.
@@ 3.1 Circuit AutoEncoder
Let $\mathcal{G}=(\mathcal{V},\mathcal{A})$ represent a circuit, where $\mathcal{V}$ denotes the set of $N$ nodes, with each node $v_{i}\in\mathcal{V}$. Following the architecture of CircuitNN (Google Deepmind, 2023; Wang et al., 2024), each node $v_{i}$ can be classified into one of three types: primary inputs (PIs), primary outputs (POs), and NAND gates, each labeled by $u_{i}\in\mathcal{U},i\in\{1,2,3\}$ respectively. The adjacency matrix $\mathcal{A}\in\{0,1\}^{N\times N}$ captures the connectivity between nodes, where $\mathcal{A}_{i,j}=1$ indicates the presence of a directed edge from $v_{i}$ to $v_{j}$.
In the circuit autoencoder framework, an encoder, denoted as $g_{E}$, encodes the circuit $\mathcal{G}$ into a latent representation $\mathbf{Z}\in\mathbb{R}^{N\times d}$ with dimensionality $d$. The encoding process for a circuit can be formulated as:
$\mathbf{Z}=g_{E}(\mathcal{V},\mathcal{A}).$ (2)
Simultaneously, a decoder $g_{D}$ aims to reconstruct the original circuit $\mathcal{G}$ from the latent representation $\mathbf{Z}$. Since node types can be directly derived from the truth table, the decoder is designed to focus on reconstructing the adjacency matrix $\mathcal{A}$, which can be formalized as follows:
$\tilde{\mathcal{G}}=(\mathcal{V},\tilde{\mathcal{A}})=(\mathcal{V},f(g_{D}(\mathbf{Z},\mathcal{V}))),$ (3)
where $\tilde{\mathcal{A}}\in\mathbb{R}^{N\times N}$ denotes the reconstructed adjacency matrix, obtained by decoding the latent representation $\mathbf{Z}$ through $g_{D}$ and applying a mapping function $f:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N\times N}$. Meanwhile, $\tilde{\mathcal{G}}$ represents the reconstructed graph. A robust encoder $g_{E}$ capable of capturing fine-grained structural information is essential to facilitate the circuit reconstruction task. We incorporate the Graphormer (Ying et al., 2021) architecture into $g_{E}$. For the decoder $g_{D}$, we adopt a simple Transformer-based (Dubey et al., 2024) architecture, as an overly powerful decoder could negatively impact the performance of the circuit tokenizer.
@@ 3.2 CircuitVQ
As mentioned in Section 3.1, we propose a circuit autoencoder architecture for the circuit reconstruction task. The outputs of $g_{E}$ and the inputs of $g_{D}$ are continuous. The circuit tokenizer is required to map the circuit to a sequence of discrete circuit tokens for masked autoregressive modeling illustrated in Section 3.3. Specifically, a circuit $\mathcal{G}$ can be tokenized to $\mathbf{Y}=[y_{1},y_{2},\cdots,y_{N}]\in\mathbb{R}^{N}$ using the circuit quantizer $\mathcal{C}$ which contains $K$ discrete codebook embeddings. Here, each token $y_{i}$ belongs to the vocabulary set $\{1,2,\dots,K\}$ of $\mathcal{C}$. Consequently, we develop a circuit tokenizer, CircuitVQ, based on the circuit autoencoder by integrating a circuit quantizer $\mathcal{C}$. As shown in Figure 2, the tokenizer comprises three components: a circuit encoder $g_{E}$, a circuit quantizer $\mathcal{C}$, and a circuit decoder $g_{D}$.
Firstly, $g_{E}$ encodes the circuit into vector representations $\mathbf{Z}$. Subsequently, $\mathcal{C}$ identifies the nearest neighbor in the codebook for $\mathbf{z}_{i}\in\mathbf{Z}$. Let $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{K}\}$ represent the codebook embeddings and $\mathbf{e}_{K}\in\mathbb{R}^{d}$. For the $i$-th node, the quantized code $y_{i}$ is determined by:
$y_{i}=\arg\min_{j}||\ell_{2}(\mathbf{z}_{i})-\ell_{2}(\mathbf{e}_{j})||_{2},$ (4)
where $j\in\{1,2,\dots,K\}$ and $\ell_{2}$ normalization is applied during the codebook lookup (Van Den Oord et al., 2017; Yu et al., 2021). This distance metric is equivalent to selecting codes based on cosine similarity. Consequently, the output of $\mathcal{C}$ for each node representation $\mathbf{z}_{i}$ the can be calculated based on the given Equation 4:
$\tilde{\mathbf{z}}_{i}$ $=\mathcal{C}(\mathbf{z}_{i})=\ell_{2}(\mathbf{e}_{y_{i}}),\text{ where }\tilde{\mathbf{z}}_{i}\in\tilde{\mathbf{Z}}.$ (5)
@@
After quantizing the circuit into discrete tokens, the $\ell_{2}$-normalized codebook embeddings $\tilde{\mathbf{Z}}=\{\tilde{\mathbf{z}}_{i}\}_{i=1}^{N}$ are fed to $g_{D}$. The output vectors $\tilde{\mathbf{X}}=\{\tilde{\mathbf{x}}_{i}\}_{i=1}^{N}=g_{D}(\tilde{\mathbf{Z}},\mathcal{V})$ are used to reconstruct the original adjacency matrix $\mathcal{A}$ of the circuit $\mathcal{G}$. Specifically, the reconstructed adjacency matrix $\tilde{\mathcal{A}}$ is derived from the output vectors $\tilde{\mathbf{X}}$ as follows:
$\tilde{\mathcal{A}}=f(\tilde{\mathbf{X}})=\sigma\left(f_{1}(\tilde{\mathbf{X}})\cdot f_{2}(\tilde{\mathbf{X}})^{T}\right),$ (6)
@@
$\mathcal{L}_{\text{rec}}=-\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\bigl{[}$ $\mathcal{A}_{ij}\log(\tilde{\mathcal{A}}_{ij})$
$+(1-\mathcal{A}_{ij})\log(1-\tilde{\mathcal{A}}_{ij})\bigr{]}.$ (7)
@@
$\mathcal{L}_{\text{vq}}=\mathcal{L}_{\text{rec}}+\|z-\text{sg}[e_{y_{i}}]\|_{2}^{2}+\beta\cdot\|\text{sg}[z]-e_{y_{i}}\|_{2}^{2}$ (8)
where $\mathrm{sg}[\cdot]$ stands for the stop-gradient operator which is an identity at the forward pass while having zero gradients during the backward pass and $\beta$ denotes the hyperparameter for commitment loss (Van Den Oord et al., 2017).
Codebook utilization. A common issue in vector quantization training is codebook collapse, where only a small proportion of codes are actively utilized. To mitigate this problem, empirical strategies (Yu et al., 2021; Jang et al., 2016) are employed. Specifically, we compute the $\ell_{2}$-normalized distance to identify the nearest code while reducing the dimensionality of the codebook embedding space (Yu et al., 2021). These low-dimensional codebook embeddings are subsequently mapped back to a higher-dimensional space before being passed to the decoder. Furthermore, we leverage the Gumbel-Softmax trick (Jang et al., 2016) to enable smoother token selection during training, ensuring that a broader range of tokens in the codebook are actively trained, thereby improving the overall utilization of the codebook.
Algorithm 1 Autoregressive Decoding of CircuitAR
0:  Masked tokens $\bar{\mathbf{Y}}=[\bar{y}_{i}]_{i=1}^{N},\forall\bar{y}_{i}=\mathit{m}$, token length $N$, total iterations $T$.
0:  Predicted tokens $\tilde{\mathbf{Y}}=[\tilde{y}_{i}]_{i=1}^{N}\forall\tilde{y}_{i}\neq\mathit{m}$.
@@
3:     Compute probabilities $p(\bar{y}_{i})\in\mathbb{R}^{K}$ for each $\bar{y}_{i}\in\bar{\mathbf{Y}}$;
4:     Initialize $\mathbf{S}\leftarrow[s_{i}]_{i=1}^{N}$, where $s_{i}=0$, and $\tilde{\mathbf{Y}}\leftarrow\bar{\mathbf{Y}}$;
@@
14:        $r\leftarrow\text{sorted}(S)[n]$; // Select the $n$-th highest score from the sorted $\mathbf{S}$ in decending order
@@
17:     $\tilde{\mathbf{Y}}\leftarrow\bar{\mathbf{Y}}$;
@@
3.3 CircuitAR
After completing the CircuitVQ training, we train CircuitAR using a graph autoregressive modeling paradigm as shown in Figure 3, where CircuitVQ functions as the tokenizer. Let $\mathbf{Y}=[y_{i}]_{i=1}^{N}$ represent the discrete latent tokens of the input circuit $\mathcal{G}$, tokenized by CircuitVQ. During the masked autoregressive training process, we sample a subset of nodes $\mathcal{V}_{s}\subset\mathcal{V}$ and replace them with a special mask token $\mathit{m}$. For the masked $\mathbf{Y}$, the latent token $\bar{y}_{i}$ is defined as:
Following Chang et al. (2022) and Li et al. (2024a), we employ a cosine mask scheduling function $\gamma(r)=\cos(0.5\pi r)$ in the sampling process. This involves uniformly sampling a ratio $r$ from the interval $[0,1]$ and then selecting $\lceil\gamma(r)\cdot N\rceil$ tokens from $\mathbf{Y}$ to mask uniformly. Let $\bar{\mathbf{Y}}=[\bar{y}_{i}]_{i=1}^{N}$ denote the output after applying the masking operation to $\mathbf{Y}$. The masked sequence $\bar{\mathbf{Y}}$ is then fed into a multi-layer transformer with bidirectional attention to predict the probabilities $p(y_{i}|\bar{\mathbf{Y}},\mathcal{T})$ for each $v_{i}\in\mathcal{V}_{s}$ under the condition of the truth table. The transformer is designed based on Llama models, each CircuitAR layer consists of a self-attention block, a cross-attention block and an FFN block. Specifically, the info of the truth table is conditioned by serving $\mathcal{T}$ as the input key and value of the cross-attention block. The training loss for CircuitAR is defined as:
$\mathcal{L}_{\text{ar}}=-\sum\limits_{\mathcal{D}}\sum\limits_{v_{i}\in\mathcal{V}_{s}}\log p(y_{i}|\bar{\mathbf{Y}},\mathcal{T}),$ (10)
where $\mathcal{D}$ represents the set of training circuits.
Autoregressive decoding. We introduce a parallel decoding method, where tokens are generated in parallel. This approach is feasible due to the bidirectional self-attention mechanism of CircuitAR. At inference time, we begin with a blank canvas $\bar{\mathbf{Y}}=[\mathit{m}]^{N}$ and the decoding process of CircuitAR follows Algorithm 1. Specifically, the decoding algorithm generate a circuit in $T$ steps. At each iteration, the model predicts all tokens simultaneously but retains only the most confident predictions following the cosine schedule (Chang et al., 2022; Li et al., 2024a). The remaining tokens are masked and re-predicted in the next iteration.
3.4 Differentiable Architecture Search
DAG Search. The reconstructed adjacency matrix $\tilde{\mathcal{A}}$ is a probability matrix that denotes the probabilities of connections between gates. However, $\tilde{\mathcal{A}}$ may contain cycles. Identifying the optimal DAG with the highest edge probabilities is an NP-hard problem, we employ a greedy algorithm to obtain a suboptimal DAG. As illustrated in Algorithm 2, the algorithm initializes $\bar{\mathcal{A}}\in\mathbb{R}^{N\times N}$ with edge probabilities and enforces basic structural rules: PIs have no indegree, POs have no outdegree, and self-loops are prohibited in circuit designs. Following this initialization, a depth-first search (DFS) is conducted to detect cycles in $\bar{\mathcal{A}}$. If no cycles are found, $\bar{\mathcal{A}}$ is a valid DAG, and the algorithm terminates. If a cycle is detected, the edge with the lowest probability within the cycle is identified and removed by setting the corresponding edge in $\bar{\mathcal{A}}$ to 0. This process repeats iteratively until no cycles remain. This greedy approach ensures the derivation of a valid DAG $\bar{\mathcal{A}}$ that approximates the optimal structure while preserving the acyclic property necessary for circuit design. The resulting DAG serves as a foundation for further refinement in the DAS process, ultimately generating a precise circuit that is logically equivalent to $\mathcal{T}$.
@@ 3.4 Differentiable Architecture Search
0:  Adjacency matrix $\tilde{\mathcal{A}}\in\mathbb{R}^{N\times N}$, PI node list $Q_{i}$, PO node list $Q_{o}$.

### v2

@@ Abstract
Logic synthesis, a critical stage in electronic design automation (EDA), optimizes gate-level circuits to minimize power consumption and area occupancy in integrated circuits (ICs). Traditional logic synthesis tools rely on human-designed heuristics, often yielding suboptimal results. Although differentiable architecture search (DAS) has shown promise in generating circuits from truth tables, it faces challenges such as high computational complexity, convergence to local optima, and extensive hyperparameter tuning. Consequently, we propose a novel approach integrating conditional generative models with DAS for circuit generation. Our approach first introduces CircuitVQ, a circuit tokenizer trained based on our Circuit AutoEncoder We then develop CircuitAR, a masked autoregressive model leveraging CircuitVQ as the tokenizer. CircuitAR can generate preliminary circuit structures from truth tables, which guide DAS in producing functionally equivalent circuits. Notably, we observe the scalability and emergent capability in generating complex circuit structures of our CircuitAR models. Extensive experiments also show the superior performance of our method. This research bridges the gap between probabilistic generative models and precise circuit generation, offering a robust solution for logic synthesis.
@@ 1 Introduction
Logic synthesis [11], a critical step in electronic design automation (EDA), transforms behavioral-level circuit designs into optimized gate-level circuits, ultimately yielding the final IC layout. The primary goal of logic synthesis is to identify the physical implementation with the fewest gates for a given circuit function. This task constitutes a challenging NP-hard combinatorial optimization problem [12]. Current logic synthesis tools [4, 32] rely on human-designed heuristics, often leading to sub-optimal outcomes.
@@
Differentiable architecture search (DAS) techniques [22, 8] offer novel perspectives on addressing challenges in this problem. Circuit functions can be represented through truth tables, which map binary inputs to their corresponding outputs. Truth tables provide a precise representation of input-output relationships, ensuring the design of functionally equivalent circuits. Inspired by this, researchers [14, 31] have begun exploring the application of DAS to synthesize circuits directly from truth tables. Specifically, [14] proposed CircuitNN, a framework that learns differentiable connection structures with logic gates, enabling the automatic generation of logic circuits from truth tables. This approach significantly reduces the complexity of traditional circuit generation. Building on this, [31] introduced T-Net, a triangle-shaped variant of CircuitNN, incorporating regularization techniques to enhance the efficiency of DAS.
@@
Despite these advancements, several challenges remain. The computational complexity of DAS grows quadratically with the number of gates, posing scalability issues. Although triangle-shaped architecture [31] partially mitigates this problem, redundancy persists. Additionally, DAS is susceptible to converging to local optima [22], where network depth and layer width require extensive searches. The challenges arise from the vast search space in DAS. Intuitively, limiting the search space through predefined parameters, including network depth, gates per layer, and connection probabilities, can significantly reduce the complexity.
@@
Recent advances [24, 1, 10, 20] in conditional generative models have demonstrated remarkable performance across language, vision, and graph generation tasks. Motivated by these developments, we propose a novel approach to circuit generation that generates preliminary circuit structures to guide DAS in generating refined circuits matching specified truth tables. Firstly, we introduce CircuitVQ, a tokenizer with a discrete codebook for circuit tokenization. Built upon our Circuit AutoEncoder framework [15, 18, 33], CircuitVQ is trained through a circuit reconstruction task. Specifically, the CircuitVQ encoder encodes input circuits into discrete tokens using a learnable codebook, while the decoder reconstructs the circuit adjacency matrix based on these tokens. Subsequently, the CircuitVQ encoder serves as a circuit tokenizer for CircuitAR pretraining, which employs a masked autoregressive modeling paradigm [7, 19]. In this process, the discrete codes function as supervision signals. After training, CircuitAR can generate discrete tokens progressively, which can be decoded into initial circuit structures by the decoder of CircuitVQ. These prior insights can guide DAS in producing refined circuits that match the target truth tables precisely. Our key contributions can be summarized as follows:
@@ 2.1 Modeling Circuit as DAG
In this work, we model the circuit as a directed acyclic graph (DAG) [5], which facilitates graph autoregressive modeling. Specifically, each node in the DAG corresponds to a logic gate, while the directed edges represent the connections between these components.
@@ 2.2 Differentiable CircuitNN
As depicted in Figure 1, CircuitNN [14] replaces traditional neural network layers with logic gates (e.g., NAND) as basic computational units, learning to synthesize circuits by optimizing logic correctness based on truth tables. During training, input connections of each gate are determined through learnable probability distributions, enabling adaptive circuit architecture modification. To enable gradient-based learning, CircuitNN transforms discrete logic operations into continuous, differentiable functions using NAND gates for simplicity. The NAND gate is logically complete, allowing the construction of any complex logic circuit. Its continuous relaxation can be defined as:
@@
Additionally, CircuitNN employs Gumbel-Softmax [16] for stochastic sampling of gate inputs. Through stochastic relaxation, gate and network outputs are no longer binary but take continuous values ranging from 0 to 1 instead. This end-to-end differentiability allows the model to learn gate input distributions using gradient descent. After training, the continuous, probabilistic circuit is converted back into a discrete logic circuit by selecting the most probable connections based on the learned probability distributions, as shown in Figure 1.
Figure 1: Illustration of differentiable CircuitNN.
@@ 3.1 Circuit AutoEncoder
Let $\mathcal{G}=(\mathcal{V},\mathcal{A})$ represent a circuit, where $\mathcal{V}$ denotes the set of $N$ nodes, with each node $v_{i}\in\mathcal{V}$. Following the architecture of CircuitNN [14, 31], each node $v_{i}$ can be classified into one of three types: primary inputs (PIs), primary outputs (POs), and NAND gates, each labeled by $u_{i}\in\mathcal{U},i\in\{1,2,3\}$ respectively. The adjacency matrix $\mathcal{A}\in\{0,1\}^{N\times N}$ captures the connectivity between nodes, where $\mathcal{A}_{i,j}=1$ indicates the presence of a directed edge from $v_{i}$ to $v_{j}$.
In the circuit autoencoder framework, an encoder, denoted as $g_{E}$, encodes the circuit $\mathcal{G}$ into a latent representation $\bm{Z}\in\mathbb{R}^{N\times d}$ with dimensionality $d$. The encoding process for a circuit can be formulated as:
$\bm{Z}=g_{E}(\mathcal{V},\mathcal{A}).$ (2)
Simultaneously, a decoder $g_{D}$ aims to reconstruct the original circuit $\mathcal{G}$ from the latent representation $\bm{Z}$. Since node types can be directly derived from the truth table, the decoder is designed to focus on reconstructing the adjacency matrix $\mathcal{A}$, which can be formalized as follows:
$\tilde{\mathcal{G}}=(\mathcal{V},\tilde{\mathcal{A}})=(\mathcal{V},f(g_{D}(\bm{Z},\mathcal{V}))),$ (3)
where $\tilde{\mathcal{A}}\in\mathbb{R}^{N\times N}$ denotes the reconstructed adjacency matrix, obtained by decoding $\bm{Z}$ through $g_{D}$ and applying a mapping function $f:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N\times N}$. Meanwhile, $\tilde{\mathcal{G}}$ represents the reconstructed graph. A robust encoder $g_{E}$ capable of capturing fine-grained structural information is essential to facilitate the circuit reconstruction task. We incorporate the Graphormer [34] architecture into $g_{E}$. For the decoder $g_{D}$, we adopt a simple Transformer-based [9] architecture, as an overly powerful decoder could negatively impact the performance of the circuit tokenizer.
@@ 3.2 CircuitVQ
As mentioned in Section 3.1, we propose a circuit autoencoder architecture for the circuit reconstruction task. The outputs of $g_{E}$ and the inputs of $g_{D}$ are continuous. The circuit tokenizer is required to map the circuit to a sequence of discrete circuit tokens for masked autoregressive modeling, illustrated in Section 3.3. Specifically, a circuit $\mathcal{G}$ can be tokenized to $\bm{Y}=[y_{1},y_{2},\cdots,y_{N}]\in\mathbb{R}^{N}$ using the circuit quantizer $\mathcal{C}$ which contains $K$ discrete codebook embeddings. Here, each token $y_{i}$ belongs to the vocabulary set $\{1,2,\dots,K\}$ of $\mathcal{C}$. Consequently, we develop a circuit tokenizer, CircuitVQ, based on the circuit autoencoder by integrating a circuit quantizer $\mathcal{C}$. As shown in Figure 2, the tokenizer comprises three components: a circuit encoder $g_{E}$, a circuit quantizer $\mathcal{C}$, and a circuit decoder $g_{D}$.
Figure 2: The training process of CircuitVQ.
Firstly, $g_{E}$ encodes the circuit into vector representations $\bm{Z}$. Subsequently, $\mathcal{C}$ identifies the nearest neighbor in the codebook for $\bm{z}_{i}\in\bm{Z}$. Let $\{\bm{e}_{1},\bm{e}_{2},\dots,\bm{e}_{K}\}$ represent the codebook embeddings and $\bm{e}_{K}\in\mathbb{R}^{d}$. For the $i$-th node, the quantized code $y_{i}$ is determined by:
$y_{i}=\arg\min_{j}||\ell_{2}(\bm{z}_{i})-\ell_{2}(\bm{e}_{j})||_{2},$ (4)
@@
where $j\in\{1,2,\dots,K\}$ and $\ell_{2}$ normalization is applied during the codebook lookup [30, 35]. This distance metric is equivalent to selecting codes based on cosine similarity. Consequently, the output of $\mathcal{C}$ for each node representation $\bm{z}_{i}$ the can be calculated based on the given Equation 4:
$\tilde{\bm{z}}_{i}$ $=\mathcal{C}(\bm{z}_{i})=\ell_{2}(\bm{e}_{y_{i}}),\text{ where }\tilde{\bm{z}}_{i}\in\tilde{\bm{Z}}.$ (5)
After quantizing the circuit into discrete tokens, the $\ell_{2}$-normalized codebook embeddings $\tilde{\bm{Z}}=\{\tilde{\bm{z}}_{i}\}_{i=1}^{N}$ are fed to $g_{D}$. The output vectors $\tilde{\bm{X}}=\{\tilde{\bm{x}}_{i}\}_{i=1}^{N}=g_{D}(\tilde{\bm{Z}},\mathcal{V})$ are used to reconstruct the original adjacency matrix $\mathcal{A}$ of the circuit $\mathcal{G}$. Specifically, the reconstructed adjacency matrix $\tilde{\mathcal{A}}$ is derived from the output vectors $\tilde{\bm{X}}$ as follows:
$\tilde{\mathcal{A}}=f(\tilde{\bm{X}})=\sigma\left(f_{1}(\tilde{\bm{X}})\cdot f_{2}(\tilde{\bm{X}})^{\top}\right),$ (6)
@@
$\mathcal{L}_{\text{rec}}=-\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\bigl{[}$ $\mathcal{A}_{ij}\log(\tilde{\mathcal{A}}_{ij})+(1-\mathcal{A}_{ij})\log(1-\tilde{\mathcal{A}}_{ij})\bigr{]}.$ (7)
@@
$\mathcal{L}_{\text{vq}}=\mathcal{L}_{\text{rec}}+\|\bm{Z}-\text{sg}[\bm{E}]\|_{2}^{2}+\beta\cdot\|\text{sg}[\bm{Z}]-\bm{E}\|_{2}^{2},$ (8)
where $\mathrm{sg}[\cdot]$ stands for the stop-gradient operator, which is an identity at the forward pass while having zero gradients during the backward pass. $\bm{E}=\{\bm{e}_{y_{i}}\}^{N}_{i=1}$ and $\beta$ denotes the hyperparameter for commitment loss [30].
Codebook utilization. A common issue in vector quantization training is codebook collapse, where only a small proportion of codes are actively utilized. To mitigate this problem, empirical strategies [35, 16] are employed. Specifically, we compute the $\ell_{2}$-normalized distance to identify the nearest code while reducing the dimensionality of the codebook embedding space [35]. These low-dimensional codebook embeddings are subsequently mapped back to a higher-dimensional space before being passed to the decoder. Furthermore, we leverage the Gumbel-Softmax trick [16] to enable smoother token selection during training, ensuring that a broader range of tokens in the codebook are actively trained, thereby improving the overall utilization of the codebook.
Algorithm 1 Autoregressive Decoding
0: Masked tokens $\bar{\bm{Y}}=[\bar{y}_{i}]_{i=1}^{N},\forall\bar{y}_{i}=\mathit{m}$, token length $N$, total iterations $T$.
0: Predicted tokens $\tilde{\bm{Y}}=[\tilde{y}_{i}]_{i=1}^{N}\forall\tilde{y}_{i}\neq\mathit{m}$.
@@
3:    Compute probabilities $p(\bar{y}_{i})\in\mathbb{R}^{K}$ for each $\bar{y}_{i}\in\bar{\bm{Y}}$;
4:    Initialize $\bm{S}\leftarrow[s_{i}]_{i=1}^{N}$, where $s_{i}=0$, and $\tilde{\bm{Y}}\leftarrow\bar{\bm{Y}}$;
@@
14:       $r\leftarrow\text{sorted}(S)[n]$; // Select the $n$-th highest score from the sorted $\bm{S}$ in decending order
@@
17:    $\tilde{\bm{Y}}\leftarrow\bar{\bm{Y}}$;
@@
0: Adjacency matrix $\tilde{\mathcal{A}}$, PI node list $Q_{i}$, PO node list $Q_{o}$.